# DB-Explore: Automated Database Exploration and Instruction Synthesis for Text-to-SQL

### v1

@@ 5.2 Ablation Study
Study on Semantic Knowledge Extraction
We conduct ablation experiments on the semantic knowledge extraction component of DB-Explore, which aligns LLMs with database domains through schema-aware semantic extraction for precise SQL generation and diversified instruction synthesis. As shown in Table 2, removing this component causes a 1.1% performance drop on the SPIDER benchmark, confirming its necessity for maintaining model efficacy.
Study on Structural Pattern Mining
As evidenced in Table 2, the structural pattern mining component is crucial to DB-Explore, as it explicitly models relational database schemas. This module enhances the LLMs’ ability to understand database topologies and structural dependencies, distinguishing it from baseline methods focused solely on direct SQL generation. Removing structural pattern mining results in a 1.9% performance drop on the SPIDER benchmark, primarily due to foreign-key errors and column selection mismatches. Notably, structural pattern mining is closely linked to empirical weighted sampling who ensures that the synthesized data better aligns with the distribution of the seed data.
Study on Progressive Instruction Synthesis
Progressive instruction synthesis enhances the complexity of instruction data by iteratively adding conditions to existing instructions. This framework effectively addresses the complexity limitations of LLM-based data synthesis. In our experiments, initial data synthesis performed well for 2-table and 3-table instructions, but for instructions involving more than three tables, LLMs exhibited a high error rate. The resulting limitations in instruction complexity significantly restrict the SQL generation capability of LLMs. As shown in Table 2, progressive instruction synthesis yielded a 0.9% performance improvement for DB-Explore. However, during the experiment, progressive instruction synthesis led to a shift in the synthesized data distribution, which resulted in suboptimal performance on the SPIDER dataset. Nonetheless, we believe that high-quality complex data can further elevate LLMs’ Text-to-SQL capabilities in real-world applications.
@@ 6 Conclusion
In this paper, we propose DB-Explore, a framework for database exploration, instruction synthesis, and fine-tuning that enhances the LLMs’ understanding of databases by extracting dynamic database knowledge, thereby unlocking the potential of LLMs in Text-to-SQL tasks. Our approach, through database multi-knowledge exploration, effectively minimizes the risks associated with insufficient database knowledge in SQL generation. We demonstrate the efficacy and superiority of our method on recent LLMs across several benchmark tests. The results show that our approach achieves satisfactory performance in Text-to-SQL with minimal computational cost. In the future, we aim to explore more database knowledge, larger LLMs, and more efficient fine-tuning frameworks to achieve robust and efficient Text-to-SQL solutions.
@@
Appendix A Appendix
Our DB-Explore framework is designed with various prompt templates aimed at enhancing GPT-4o’s performance in tasks such as semantic knowledge extraction, structural pattern mining, and progressive instruction synthesis. We generated a total of 6,497 question-SQL pairs on the SPIDER dataset and 6,687 question-SQL pairs on the BIRD dataset. In this appendix, we provide the prompts of Database Multi-Knowledge Exploration used in our pipeline as shown in Figure 5 to 7.
Figure 5: The prompt of Semantic Knowledge Extraction.
Figure 6: The prompt of Structural Pattern Mining.
Figure 7: The prompt of Progressive Instruction Synthesis.

### v2

@@ 5.2 Ablation Study
Study on Semantic Knowledge Extraction and Structural Pattern Mining We conducted ablation experiments to quantify the individual and joint effects of Semantic Knowledge Extraction and Structural Pattern Mining on DB-Explore’s performance. As shown in Table 3, when the Semantic Knowledge Extraction module is removed, accuracy decreases by 1.1% on SPIDER and 0.4% on BIRD, indicating its role in extracting entity categories and predicate relationships that guide precise column and value mapping. By contrast, disabling Structural Pattern Mining which captures foreign-key dependencies and the global topology of the database schema results in a much larger 3.3% drop on SPIDER and 2.7% drop on BIRD, primarily due to increased join-path selection failures and ambiguous column predictions. Due to the inherent semantic information present in the synthesized data, Structural Pattern Mining brings even larger gains than Semantic Knowledge Extraction.
Study on Progressive Instruction Synthesis To assess the impact of gradually increasing query complexity, we evaluated DB-Explore with and without the Progressive Instruction Synthesis module. Incorporation of this module yields a 2.4% improvement on SPIDER and a 3.0% gain on BIRD, with the majority of benefits observed on moderate and challenging queries that require multi-table joins and nested filters.
Study on Instruction Filtering Finally, we examined the necessity of our Instruction Filtering component, which prunes hallucinated or incomplete instructions produced during synthesis. Without this filter, performance on SPIDER declines by 3.5% and on BIRD by 3.8%, reflecting the disruptive influence of low-quality examples. This ablation confirms that rigorous filtering is indispensable for maintaining instruction quality and preventing performance degradation caused by spurious or malformed queries.
@@ 6 Conclusion
In this paper, we propose DB-Explore, a framework for database exploration, instruction synthesis, and fine-tuning that enhances the LLMs’ understanding of databases by extracting dynamic database knowledge, thereby unlocking the potential of LLMs in Text-to-SQL tasks. Our approach, through database multi-knowledge exploration, effectively minimizes the risks associated with insufficient database knowledge in SQL generation. We demonstrate the efficacy and superiority of our method on recent LLMs across several benchmark tests. The results show that our approach achieves state-of-the-art performance in Text-to-SQL with minimal computational cost. In the future, we aim to explore more database knowledge, larger LLMs, and more efficient fine-tuning frameworks to achieve robust and efficient Text-to-SQL solutions.